Hi. I have an exam tomorrow and I am working on a practice test and I have a question about a problem.
Put each equation into slope-intercept form: y = mx + b.
2x + y = 2
So I got the answer: y = 2x + 2
The answer key says: y = -2x + 2
How did it become a negative -2x? How would I know when doing a problem that something with no negatives become a negative in the answer.

Found 2 solutions by jigs_ipan, feliz1965:
Answer by jigs_ipan(5)

that works like this:
2x+y=2
to transform it in slope intercept form:
(2x+y)-(2x)=2-(2x) (subtract both sides by 2x so that the only variable in the left is "y")
so the result is:
2x+y-2x=2-2x
y=2-2x or y=-2x+2

Answer by feliz1965(151)

You did very well. The reason for -2x is basically because when crossing the equal sign, in either direction, the sign of the term, variable or constant must be changed to the opposite.
For example, solve for m:
3m + 5 = -2m
If I move 3m to the right side of the equation, it becomes -3m.
If I decide to move -2m to the left side of the equation, -2m becomes 2m.
The same applies for constants.
Is this clear?
Feliz